OpenStudy (jmartinez638):

Help me with these statements: 1)The definite integral of a polynomial function of odd degree over [-3,3] will __________ be equal to 0. 2)A function f that is continuous on [a,b] and has zeroes at a and b will __________ have a horizontal tangency at some x in (a,b).

OpenStudy (jmartinez638):

These are Always, Sometimes, Never problems.

OpenStudy (jmartinez638):

If c is a critical number of f and f”(c) = 2, then f _____________ has a point of inflection at x = c.

OpenStudy (solomonzelman):

Think about it... \(\large\color{#000000 }{ \displaystyle F(x)=\int_{-a}^{a} x^{2n+1}dx=\frac{1}{2n+2}x^{2n+2} }\) And then, F(-3)=F(3), because of the symmetry of even power. So, F(3) - F(-3) = 0 HOWEVER, some terms can be \(\large\color{#000000 }{ \displaystyle f(x)=\int_{-a}^{a} x^{2n}dx=\frac{1}{2n+1}x^{2n+1} }\) and this is not necessarily 0

OpenStudy (solomonzelman):

I am saying that sometimes it can be a poly with degree 2n+1, and another lesser term x^2n...

OpenStudy (solomonzelman):

there is a test for inflection

OpenStudy (solomonzelman):

let's take some sufficiently small b, let's say b=0.06 So, if f"(c+b) and f"(c-b) have different signs, then concavity indeed changes, and the point x=c is indeed inflection

OpenStudy (jmartinez638):

Yes.

OpenStudy (solomonzelman):

However, if f"(c+b) and f"(c-b) have same signs then x=c is NOT an inflection.

OpenStudy (solomonzelman):

So your answer would be?

OpenStudy (jmartinez638):

Sometimes

OpenStudy (solomonzelman):

oh sorry I read f”(c) = 0

OpenStudy (solomonzelman):

if =0 , then c is possibly an inflcetion.

OpenStudy (solomonzelman):

but f”(c) = 2 never an inflcetion at x=c

OpenStudy (jmartinez638):

O ok

OpenStudy (solomonzelman):

I got to go... (f is concave up at c if f''(c)=2)

OpenStudy (jmartinez638):

If f(x) < 0 for all \[x \in [a,b], \] then the area between f(x) and x=0 can _____ be accurately calculated with \[\int\limits_{b}^{a} f(x) dx.\]